# Accelerating Tactile Internet with QUIC: An Exploration of its Security and Privacy Attacks

### v1

@@
Abstract
The Tactile Internet paradigm is set to revolutionize human society by enabling skill-set delivery and haptic communication over ultra-reliable, low-latency networks. The emerging sixth-generation (6G) mobile communication systems are envisioned to underpin this Tactile Internet ecosystem at the network edge by providing ubiquitous global connectivity. However, apart from a multitude of opportunities of the Tactile Internet, security and privacy challenges emerge at the forefront. We believe that the recently standardized QUIC protocol, characterized by end-to-end encryption and reduced round-trip delay would serve as the backbone of Tactile Internet. In this article, we envision a futuristic scenario where a QUIC-enabled network uses the underlying 6G communication infrastructure to achieve the requirements for Tactile Internet. Interestingly this requires a deeper investigation of a wide range of security and privacy challenges in QUIC, that need to be mitigated for its adoption in Tactile Internet. Henceforth, this article reviews the existing security and privacy attacks in QUIC and their implication on users. Followed by that, we discuss state-of-the-art attack mitigation strategies and investigate some of their drawbacks with possible directions for future work.
I Introduction
The Tactile Internet [1], driven by recent technological advancements, focuses on transmitting touch in real-time via suitable haptic equipment and robotics over a robust communication network. Tactile Internet promises to deliver real-time control through ultra-responsive and ultra-reliable network connectivity over a distance to create a more sustainable ecosystem. This is expected to revolutionize the entire human society by shifting the focus from mere content delivery to a new system of skill-set delivery over the Internet. However, its implementation at the infrastructure level demands ensuring low latency with highly reliable communication and privacy coupled with a very high level of security. This is because Tactile Internet is designed to serve highly critical applications where human senses (touch, video, and audio) and muscular movements directly interact with machines, i.e., robots. Hence, cybersecurity and quick reaction time ($<1$ ms) are two primary requirements of the Tactile Internet [1].
Any significant delay can induce cyber-sickness in users due to unsynchronized senses, occurring as a result of a time lag ($>1$ ms) between visual and tactile movements. Achieving global connectivity with a high data rate and minimizing jitter are additional challenges. Given such non-elastic requirements, the emerging sixth-generation (6G) communication system, offering ubiquitous connectivity, is poised to serve as the Tactile Internet’s network backbone. While existing 5G systems meet the throughput requirements, the integration of terahertz (THz) communication and millimeter-wave (mmWave) in 6G ensures higher data rates and ultra-low-latency (0.1 ms delay), making it ideal for the Tactile Internet [2].
Figure 1: A high-level architectural view of a QUIC-enabled futuristic Tactile Internet infrastructure using 6G as the underlying mobile communication network.
It is evident from the above discussion that security/privacy breaches and packet losses in the Tactile Internet environment can lead to detrimental effects, particularly if it is used in mission-critical applications. In this regard, the traditional transport layer protocols such as TCP and UDP are inadequate as they are not designed to ensure security. Moreover, the “TCP head-of-line blocking” problem  [3] poses a significant bottleneck for Tactile Internet applications, as packet loss from one source leads to buffering of successive packets from all sources (e.g., different sensors). Given the high redundancy in sensory data, receiving every generated packet is not mandatory. Hence, the Internet Engineering Task Force’s (IETF) recently standardized transport protocol, QUIC (see: §II), is a suitable alternative.
QUIC is a connection-oriented, end-to-end encrypted protocol based on UDP and built on top of TLS 1.3, thus supporting $0-RTT$. QUIC allows sending parallel and independent data streams which are logically separate from one another. This ensures fast and reliable in-order delivery of independent data streams, thereby avoiding the head-of-line blocking problem faced by TCP [3]. Though QUIC was primarily designed to handle Web traffic, its unique features (low-latency delivery, security, $0-RTT$ performance benefits) are suitable for interactive Tactile Internet applications that need to transmit kinesthetic and tactile data collected from various sensors with potential low delay and high security. Moreover, QUIC is already under the deployment stage across multiple application scenarios such as DNS, VPNs and beyond other than just the Web [3]. Inspired by these facts, we envision QUIC as a potential candidate for 6G-based Tactile Internet applications as depicted in Fig. 1. Consequently, a deeper investigation of the various security (see: §II) and privacy (see: §III) challenges in QUIC is needed.
This article provides a comprehensive overview of the security and privacy issues in QUIC. We believe this will help practitioners and system designers take appropriate measures before deploying QUIC in a Tactile Internet environment. We also discuss a few mitigation strategies and highlight some of their drawbacks (see: §IV), providing guidance for the research community to delve deeper and propose alternative or improved solutions to the challenges outlined in this article.
Figure 2: QUIC connection establishment with attack vectors. The packets or tuples under attack are represented in blue, whereas the attacks are highlighted in red.
II Security Attacks on QUIC
@@
We begin this section by providing a deep dive into QUIC handshakes, as several security attacks exploit the packets exchanged during QUIC connection establishment. We then discuss the relevant security attacks on QUIC (see: Table I) applicable to a Tactile Internet environment.
@@
QUIC employs end-to-end encryption, necessitating agreement on multiple parameters between endpoints (client and server). This agreement occurs through a TLS 1.3 handshake, as depicted in Fig. 2. A typical QUIC segment’s header fields are also illustrated. When a client lacks prior knowledge of a server, it initiates the process by sending an Initial Client Hello (CHLO) message to the server. Upon receiving CHLO, the server responds with a reject reject (REJ) packet to the client, containing tuples: (i) server config (scfg) with Diffie-Hellman public value, (ii) source-address token (stk) comprising timestamp from the server and client’s IP address, (iii) server authentication certificate chain, and (iv) server config signature. After receiving the server config, the client authenticates it using the certificate chain and the server signature. The client then creates a Complete CHLO, including its ephemeral Diffie-Hellman key, and sends it back to the server. With the completion of the handshake, the client obtains the initial keys for communication, enabling it to send an encrypted connection request to the server. The server responds with a server hello (SHLO) packet, encrypted with initial keys and the server’s ephemeral keys. Both client and server then calculate their forward secure keys to be used for subsequent communication rounds. The client can now engage in a 00-RTT handshake, bypassing the initial handshake mechanism, by sending a Complete CHLO.
II-A 0-RTT Attack
@@
During the initiation of a connection setup between the client and server, a 00-RTT attack can occur when the attacker intercepts the unencrypted initial CHLO packet and spoofs it (Fig. 2). This opens the door to two potential attacks [4]:
@@
In this regard, Cao et al. [4] have detailed the core concepts and implementation of the 00-RTT attack, emphasizing its potential to result in denial-of-service. They have introduced a mathematical attack description model based on finite state machines to demonstrate the QUIC protocol’s vulnerability, showcasing the attack process and confirming QUIC’s susceptibility to the described attacks.
II-B Replay Attack
A replay attack occurs when an intruder intercepts a legitimate network transmission and later resends it to deceive the system into treating the re-transmitted data as authentic. Lychev et al. [5] conducted two types of replay attacks (Fig. 2) on the QUIC implementation in Chromium:
@@ II-B Replay Attack
Source Address Token Replay Attack: In this attack, the intruder replays the source-address token (stk) of a client to the server that issued it, allowing the creation of multiple additional connections. This will make the server establish both initial keys and final forward-secure keys. Although subsequent handshake steps would fail, the attacker could potentially launch a denial-of-service (DoS) attack on the server by generating numerous connections on behalf of various clients, leading to a depletion of computational and memory resources.
Fischlin et al. [6] addressed replay attacks in the 00-RTT QUIC handshake. While 00-RTT uses registers for storing nonces to prevent repetition, in QUIC, once the server rejects the 00-RTT message, the attacker resends the rejected message using a second key to ensure proper delivery. This results in duplicate processing of the same data by the server.
TABLE I: Overview of various Security and Privacy Attacks and their Effects on QUIC
@@
Connection Failure; Server Load [5]
@@
[5]
Request Forgery Attack Traffic Amplification [7]
@@
[10]
Website Fingerprinting User Privacy breach [11, 12]
II-C Manipulation Attack
A manipulation attack aims to compromise the key agreement process, leading the client and server to agree on distinct keys. It is achieved by altering unprotected packet fields, such as connection id (cid) or the source-address token (stk) (Fig. 2) that are utilized as input to the key derivation process. Lychev et al. [5] describe the following two types of manipulation attacks in QUIC:
@@ II-C Manipulation Attack
II-D Request Forgery Attack
This attack involves instigating a client to send deceptive requests to another client. In [7], Gbur et al. assume a scenario where the attacker has complete control over packets delivered to the victim, enabling three types of forgery attacks:
Server Initial Request Forgery (SIRF): The attacker initiates a QUIC handshake with the server, where the packet’s source IP address and port fields are spoofed. This tricks the victim to assume that the connection is from a genuine server.
@@ II-D Request Forgery Attack
II-E State-Overflow and Resource Exhaustion Attacks
@@
A study [8] evaluates these attacks on real-world QUIC traffic based on active measurements using the UCSD network telescope. Results show that, similar to TCP SYN floods, the QUIC handshake is susceptible to resource exhaustion attacks. Notably, 98% of these attacks target QUIC servers of well-known companies, with Google being the primary target (58%) and Facebook accounting for 25% of the attacks. While QUIC originally supports RETRY messages to mitigate resource exhaustion attacks, their implementation adds an extra round-trip time (RTT), conflicting with QUIC’s claimed performance gains. This happens as a RETRY message typically precedes a QUIC handshake and forces the client to prove its authenticity by responding with a unique token. The study [8] emphasizes that during experimentation, no RETRY messages were observed, indicating a lack of practical defense mechanisms in QUIC deployments.
II-F QUIC Flood DDoS Attack
As previously explained, a QUIC flood DDoS attack occurs when an attacker overwhelms a QUIC server with a large volume of data, causing the victim server to slow down significantly or crash. Defending against DDoS attacks in QUIC is challenging because QUIC is based on UDP, which provides minimal or no information for blocking illegitimate traffic. Additionally, as QUIC packets are encrypted, the victim server cannot easily verify the legitimacy of the data source. While some Content Distribution Networks (CDNs) like Cloudflare have been successful in mitigating QUIC floods [9], a study in [8] indicates that these attacks persist in the real world. In [8], the authors reported that the Internet is vulnerable to four QUIC flood attacks every hour. Among these, 51% of the attacks occur concurrently with TCP/ICMP floods, while another 40% target the victim sequentially. However, QUIC floods have a shorter duration, lasting only $255\ seconds$, compared to their TCP/ICMP counterparts. However, such shorter duration may lead to faster resource exhaustion or system breakdown, presenting an aspect that is yet to be explored.
III Privacy Attacks on QUIC
@@
III-A User Tracking via QUIC
@@
The study [10] reveals QUIC’s privacy violation, enabling trackers to map users based on tokens provided during a 00-RTT connection setup attempt. Analysis of popular browsers (Chrome) revealed insufficient protective measures against Web user tracking. Such monitoring was particularly effective in resource-constrained scenarios due to lower bandwidth requirements and delays compared to HTTP cookies or traditional browser fingerprinting.
In a use-case scenario (Fig. 3), despite using encrypted QUIC connection, an attacker observing traffic traces during a real-time haptic signal transfer session (such as remote music training), can still gather information about students and the teacher. This information could be strategically exploited for user profiling and collecting sensitive details like gender and location.
Figure 3: A use-case scenario of Online Education where a network attacker can eavesdrop on the QUIC connection to launch privacy attacks.
III-B Website Fingerprinting
@@
An encrypted transmission protocol is susceptible to website fingerprinting (WFP) attacks if adversaries can deduce a user’s visited websites by monitoring the transmission channel. As QUIC encrypts data, adversaries target unencrypted handshake packets. Further, as QUIC is based on a request-response model, it allows attackers to distinguish traffic from different Web resources due to the relatively fixed browser rendering sequence [11]. This weakens privacy benefits achieved through encryption under WFP attacks.
@@
Zhan et al. [11] explored WFP attack vulnerabilities in both QUIC variants: (i) GQUIC (Google’s QUIC) and (ii) IQUIC (IETF’s QUIC). GQUIC is more vulnerable in restricted traffic scenarios, while both protocols exhibit similar vulnerabilities under typical traffic. Multiple WFP attacks with only 40 packets achieved 95.4% and 95.5% accuracy for GQUIC and IQUIC, respectively.
@@
To counter WFP attacks, IQUIC introduces a PADDING frame, but studies suggest network-level padding inefficiency against adversaries capable of observing traffic traces. Current research [12] focuses on integrating defenses directly into client applications (e.g., browsers) by effectively using the PADDING frame to thwart adversary’s attempt of traffic analysis in Virtual Private Networks (VPNs).
@@
A summary of various security and privacy attacks on QUIC is presented in Table I.
IV Potential Mitigation Strategies
In QUIC, connection establishment between the endpoints is achieved through a TLS 1.3 handshake where the client sends an initial ClientHello to the server over a UDP socket. However, this introduces several security and privacy challenges as sensitive parameters are negotiated between the endpoints during such TLS connections. Though TLS 1.3 encrypts most of the information exchanged, an on-path attacker can still learn certain sensitive private information such as Server Name Identification (SNI) extension (see: § II) available in the ClientHello message. It is also evident from § II that most of the attacks occur due to the unencrypted packet exchange during connection establishment as eavesdroppers get access to a plethora of metadata during this phase.

### v3

# How Resilient is QUIC to Security and Privacy Attacks?

@@
Abstract.
QUIC has rapidly evolved into a cornerstone transport protocol for secure, low-latency communications, yet its deployment continues to expose critical security and privacy vulnerabilities, particularly during connection establishment phases and via traffic analysis. This paper systematically revisits a comprehensive set of attacks on QUIC and emerging privacy threats. Building upon these observations, we critically analyze recent IETF mitigation efforts, including TLS Encrypted Client Hello (ECH), Oblivious HTTP (OHTTP) and MASQUE. We analyze how these mechanisms enhance privacy while introducing new operational risks, particularly under adversarial load. Additionally, we discuss emerging challenges posed by post-quantum cryptographic (PQC) handshakes, including handshake expansion and metadata leakage risks. Our analysis highlights ongoing gaps between theoretical defenses and practical deployments, and proposes new research directions focused on adaptive privacy mechanisms. Building on these insights, we propose future directions to ensure long-term security of QUIC and aim to guide its evolution as a robust, privacy-preserving, and resilient transport foundation for the next-generation Internet.
copyright: none
1. Introduction
QUIC (Iyengar and Thomson, 2021; Duke, 2023) is a connection-oriented end-to-end encrypted transport protocol based on UDP and built on top of TLS 1.3. Designed to improve upon TCP+TLS by integrating transport and cryptographic handshake layers, QUIC offers features such as reduced connection establishment latency, encrypted transport metadata, and improved resilience to network disruptions. QUIC allows sending parallel and independent data streams which are logically separate from one another, thus ensuring fast and reliable in-order delivery of independent data streams, which avoids the head-of-line blocking problem that TCP typically faces (Chiariotti et al., 2021).
Despite these advancements, QUIC’s real-world deployments continue to expose critical security and privacy vulnerabilities. Although QUIC was designed primarily to handle Web traffic, its unique features (such as low-latency, $0-RTT$ benefits) and flexibility (multi-streaming) are suitable for a wide range of applications such as DNS, VPNs and beyond, other than just the Web (HTTP traffic) (Chiariotti et al., 2021). As QUIC’s adoption accelerates across web services, content delivery networks, and mobile platforms, systematically understanding and addressing the various security (see: §2) and privacy (see: §3) challenges becomes vital-both for protocol designers and for ongoing standardization efforts within the IETF.
Several recent studies have provided valuable overviews of QUIC’s threat landscape. The study (Joarder and Fung, 2024) conducted a comprehensive survey of QUIC security and privacy vulnerabilities, threats, and attacks, outlining broad threat categories. Their work proposes high-level research directions but does not deeply examine mitigation efforts underway within the IETF. Meanwhile, Chatzoglou et al. (Chatzoglou et al., 2023) revisited QUIC’s attack vectors through a detailed review and a hands-on evaluation, focusing on the feasibility of various attacks against early QUIC deployments. However, this study also does not address the implications of newly emerging standards or the operational challenges posed by post-quantum cryptographic integration into QUIC. Thus, while these works underscore the importance of securing QUIC, critical gaps remain. In particular, there is a lack of integrated analysis connecting recent privacy-enhancing mechanisms, their impact and the challenges introduced. Moreover, how new mitigation strategies themselves could create operational vulnerabilities — such as new traffic fingerprinting opportunities — remains underexplored.
In this paper, we address these gaps by systematically revisiting modern attacks on QUIC and its privacy issues. We critically evaluate the security and operational impacts of the emerging IETF mitigation efforts (see: §4), identifying risks, and outlining future research directions. Our analysis highlights how stronger privacy protections can inadvertently introduce new operational risks under adversarial conditions, motivating the need for resilience enhancements. We also examine the challenges posed by post-quantum cryptographic integration into QUIC (see: §5). Through this work, our goal is to contribute a timely and actionable analysis that supports both academic research and practical engineering efforts within the IETF community toward building a more secure, privacy-preserving QUIC transport layer.
2. Security Attacks on QUIC
@@
We begin this section by providing a deep dive into QUIC handshakes, as several security attacks exploit the packets exchanged during QUIC connection establishment. We then discuss the relevant security attacks on QUIC (see: Table 1).
Figure 1. QUIC connection establishment with attack vectors. The packets or tuples under attack are represented in blue, whereas the attacks are highlighted in red.
@@
QUIC employs end-to-end encryption, necessitating agreement on multiple parameters between endpoints (client and server). This agreement occurs through a TLS 1.3 handshake, as depicted in Fig. 1. A typical QUIC segment’s header fields are also illustrated. When a client lacks prior knowledge of a server, it initiates the process by sending an Initial Client Hello (CHLO) message to the server. Upon receiving CHLO, the server responds with a reject (REJ) packet to the client, containing tuples: (i) server config (scfg) with Diffie-Hellman public value, (ii) source-address token (stk) comprising timestamp from the server and client’s IP address, (iii) server authentication certificate chain, and (iv) server config signature. After receiving the server config, the client authenticates it using the certificate chain and the server signature. The client then creates a Complete CHLO, including its ephemeral Diffie-Hellman key, and sends it back to the server. With the completion of the handshake, the client obtains the initial keys for communication, enabling it to send an encrypted connection request to the server. The server responds with a server hello (SHLO) packet, encrypted with initial keys and the server’s ephemeral keys. Both client and server then calculate their forward secure keys to be used for subsequent communication rounds. The client can now engage in a 00-RTT handshake, bypassing the initial handshake mechanism, by sending a Complete CHLO.
2.1. 0-RTT Attack
@@
During the initiation of a connection setup between the client and server, a 00-RTT attack can occur when the attacker intercepts the unencrypted initial CHLO packet and spoofs it (Fig. 1). This opens the door to two potential attacks (Cao et al., 2019):
@@
In this regard, Cao et al. (Cao et al., 2019) have detailed the core concepts and implementation of the 00-RTT attack, emphasizing its potential to result in denial-of-service. They have introduced a mathematical attack description model based on finite state machines to demonstrate the QUIC protocol’s vulnerability, showcasing the attack process and confirming QUIC’s susceptibility to the described attacks.
2.2. Replay Attack
A replay attack occurs when an intruder intercepts a legitimate network transmission and later resends it to deceive the system into treating the re-transmitted data as authentic. Lychev et al. (Lychev et al., 2015) conducted two types of replay attacks (Fig. 1) on the QUIC implementation in Chromium:
@@ 2.2. Replay Attack
Source Address Token Replay Attack: An intruder replays the source-address token (stk) of a client to the issuing server, allowing the creation of multiple additional connections. This will make the server establish both initial keys and final forward-secure keys. Although subsequent handshake steps would fail, the attacker could potentially launch a denial-of-service (DoS) attack on the server by generating numerous connections on behalf of various clients, leading to a depletion of computational and memory resources.
Fischlin et al. (Fischlin and Günther, 2017) addressed replay attacks in the 00-RTT QUIC handshake. While 00-RTT uses registers for storing nonces to prevent repetition, in QUIC, once the server rejects the 00-RTT message, the attacker resends the rejected message using a second key to ensure proper delivery. This results in duplicate processing of the same data by the server.
Table 1. Overview of various Security and Privacy Attacks and their Effects on QUIC
@@
Connection Failure; Server Load (Lychev et al., 2015)
@@
(Lychev et al., 2015)
Request Forgery Attack Traffic Amplification (Gbur and Florian, 2023)
@@
(Sy et al., 2019)
Website Fingerprinting User Privacy breach (Zhan et al., 2021; Smith et al., 2022)
2.3. Manipulation Attack
A manipulation attack aims to compromise the key agreement process, leading the client and server to agree on distinct keys. It is achieved by altering unprotected packet fields, such as connection id (cid) or the source-address token (stk) (Fig. 1) that are utilized as input to the key derivation process. Lychev et al. (Lychev et al., 2015) describe the following two types of manipulation attacks in QUIC:
@@ 2.3. Manipulation Attack
2.4. UDP Hole Punching Bypass Attack
Another attack scenario that may arise due to QUIC’s inherent characteristics is the UDP hole punching bypass attack. QUIC, commonly used on top of UDP, is known to be compatible with stateful firewalls. UDP typically require an examination of individual packets to establish a connection. However, stateful firewalls do not provide this level of verification for UDP connections, which leaves them vulnerable. In (Gbur and Tschorsch, 2021), Konrad et al. highlights that a server can be compromised by an attacker who uses remote code execution. Owing to the connectionless feature of UDP, any 5-tuple message from the server side may be captured to create a hole for the consequent packets with similar 5 tuples. This allows the attacker to keep the connection active by sending packets at regular intervals. As a result, the connection may remain open for a significant period of time, even after it has been terminated.
2.5. Request Forgery Attack
This attack involves instigating a client to send deceptive requests to another client. In (Gbur and Florian, 2023), Gbur et al. assume a scenario where the attacker has complete control over packets delivered to the victim, enabling three types of forgery attacks:
Server Initial Request Forgery (SIRF): An attacker initiates a QUIC handshake with the server, by spoofing the packet’s source IP address and port fields. This tricks the victim to assume that the connection is from a genuine server.
@@ 2.5. Request Forgery Attack
2.6. State-Overflow and Resource Exhaustion Attacks
@@
A study (Nawrocki et al., 2021) evaluates these attacks on real-world QUIC traffic based on active measurements using the UCSD network telescope. Results show that, similar to TCP SYN floods, the QUIC handshake is susceptible to resource exhaustion attacks. Notably, 98% of these attacks target QUIC servers of well-known companies, with Google being the primary target (58%) and Facebook accounting for 25% of the attacks. While QUIC originally supports RETRY messages to mitigate resource exhaustion attacks, their implementation adds an extra round-trip time (RTT), conflicting with QUIC’s claimed performance gains. This happens as a RETRY message typically precedes a QUIC handshake and forces the client to prove its authenticity by responding with a unique token. The study (Nawrocki et al., 2021) emphasizes that during experimentation, no RETRY messages were observed, indicating a lack of practical defense mechanisms in QUIC deployments.
In order to limit the effect of resource exhaustion attacks, QUIC originally supports RETRY messages. A RETRY message typically precedes a QUIC handshake and forces the client to prove its authenticity by responding with a unique token. However, this adds another RTT which eventually conflicts with the performance gains claimed by QUIC. In this regard, the study (Nawrocki et al., 2021) further reveals that they did not observe any RETRY messages during their experimentation which reconfirms the lack of practical defense mechanisms in QUIC deployments.
2.7. QUIC Flood DDoS Attack
As previously explained, a QUIC flood DDoS attack occurs when an attacker overwhelms a QUIC server with a large volume of data, causing the victim server to slow down significantly or crash. Defending against DDoS attacks in QUIC is challenging because QUIC is based on UDP, which provides minimal or no information for blocking illegitimate traffic. Additionally, as QUIC packets are encrypted, the victim server cannot easily verify the legitimacy of the data source. While some Content Distribution Networks (CDNs) like Cloudflare have been successful in mitigating QUIC floods (Inc., 2023), a study in (Nawrocki et al., 2021) indicates that these attacks persist in the real world. In (Nawrocki et al., 2021), the authors reported that the Internet is vulnerable to four QUIC flood attacks every hour. Among these, 51% of the attacks occur concurrently with TCP/ICMP floods, while another 40% target the victim sequentially. However, QUIC floods have a shorter duration, lasting only $255\ seconds$, compared to their TCP/ICMP counterparts. However, such shorter duration may lead to faster resource exhaustion or system breakdown, presenting an aspect that is yet to be explored.
3. Privacy Attacks on QUIC
@@
3.1. User Tracking via QUIC
@@
The study (Sy et al., 2019) reveals QUIC’s privacy violation, enabling trackers to map users based on tokens provided during a 00-RTT connection setup attempt. Analysis of popular browsers (Chrome) revealed insufficient protective measures against Web user tracking. Such monitoring was particularly effective in resource-constrained scenarios due to lower bandwidth requirements and delays compared to HTTP cookies or traditional browser fingerprinting.
3.2. Website Fingerprinting
@@
An encrypted transmission protocol is susceptible to website fingerprinting (WFP) attacks if adversaries can deduce a user’s visited websites by monitoring the transmission channel. As QUIC encrypts data, adversaries target unencrypted handshake packets. Further, as QUIC is based on a request-response model, it allows attackers to distinguish traffic from different Web resources due to the relatively fixed browser rendering sequence (Zhan et al., 2021). This weakens privacy benefits achieved through encryption under WFP attacks.
@@
Zhan et al. (Zhan et al., 2021) explored WFP attack vulnerabilities in both QUIC variants: (i) GQUIC (Google’s QUIC) and (ii) IQUIC (IETF’s QUIC). GQUIC is more vulnerable in restricted traffic scenarios, while both protocols exhibit similar vulnerabilities under typical traffic. Multiple WFP attacks with only 40 packets achieved 95.4% and 95.5% accuracy for GQUIC and IQUIC, respectively.
@@
To counter WFP attacks, IQUIC introduces a PADDING frame, but studies suggest network-level padding inefficiency against adversaries capable of observing traffic traces. Current research (Smith et al., 2022) focuses on integrating defenses directly into client applications (e.g., browsers) by effectively using the PADDING frame to thwart adversary’s attempt of traffic analysis in Virtual Private Networks (VPNs).
@@
A summary of various security and privacy attacks on QUIC is presented in Table 1.
4. Potential Mitigation Strategies
The previous two sections have broadly highlighted the emerging security and privacy challenges, particularly centered around metadata leakage and handshake vulnerabilities. As QUIC matures into the de facto transport protocol for modern secure communications, addressing its evolving security and privacy vulnerabilities has become a critical research and operational priority. This section critically evaluates emerging mitigation strategies standardized or proposed within the IETF, outlines unresolved challenges, and identifies promising research directions to make QUIC more resilient for future Internet deployments.
4.1. Encrypting Handshake Metadata